1. A pair of fair dice is cast. Calculate the probability that the two dice do not show the same number

a. 0.1389

b. 0.0278

c. 0.1667

d. 0.8333


1
Expert's answer
2021-01-25T10:14:05-0500

Total sample space = 36

Sample outcome for same number = {(1,1), (2,2), (3,3), (4,4), (5,5), (6,6)}

P(dice don't show same number) = 1 - P(same number) = "1 - \frac{6}{36} = \frac{5}{6} = 0.8333"

Answer: d. 0.8333